OpenStudy (anonymous):

find the integers,given the following sum is 6 product is -27

OpenStudy (anonymous):

9, -3

OpenStudy (anonymous):

eq.1 x+y=6 x=6-y (eq1' from eq 1) eq3 xy=-27 substitute eq1' to eq 3 (6-y)y=-27 \[6y-y ^{2}=-27\] \[y ^{2}-6y-27=0\] factor (y-9)(y+3)=0 y-9=0 y=9 substitute y=9 to eq1 x+9=6 x=6-9 x=-3

OpenStudy (anonymous):

do you know how to do it now????